# TRADES: Generating Realistic Market Simulations with Diffusion Models

### v2

@@ 4.1. Generation with TRADES
(4) $\boldsymbol{\epsilon}_{\theta}:\left(\mathbf{x}^{g}_{t}\in\mathbb{R}^{S\times K},\>\mathbf{x}^{c}_{0}\in\mathbb{R}^{M\times K},\>t\in\mathbb{R}^{K}\right)\rightarrow\boldsymbol{\epsilon}_{t}\in\mathbb{R}^{S\times K}$
where $M+S=N$. We set $M=N-1$ and $S=1$ for our experiments. Using $S>1$ increases the efficiency, but also the task complexity because, in a single generative step, we generate $S$ samples. We exploit the parametrization proposed in (Ho et al., 2020) described in Eq. (5) to estimate the mean term.
[…]
We do not rely on a fixed schedule as in (Ho et al., 2020) regarding the variance term. Inspired by (Nichol and Dhariwal, 2021), we learn it as in Eq. (6).
(6) $\mathbf{\Sigma}_{\theta}(\mathbf{x}_{t}^{g},\mathbf{x}_{0}^{c},t)=\exp({v\log\beta_{t}+(1-v)\log\tilde{\beta}_{t}}),$
where $v$ is the neural network output, together with $\epsilon_{\theta}$. Nichol and Dhariwal (Nichol and Dhariwal, 2021) found that this choice improves the negative log-likelihood, which we try to minimize. After computing $\boldsymbol{\epsilon}_{t}$ and $\boldsymbol{\sigma}_{t}$, we denoise $\mathbf{x}_{t-1}^{g}$ as in Eq (7).
(7) $\mathbf{x}^{g}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}^{g}_{t}-\frac{\beta_{t}}{\sqrt{1-\mkern 2.0mu\overline{\mkern-2.0mu\alpha\mkern-2.0mu}\mkern 2.0mu_{t}}}\boldsymbol{\epsilon}_{\theta}(\mathbf{x}^{g}_{t},\mathbf{x}^{c}_{0},t)\right)+\boldsymbol{\sigma}_{t}\mathbf{z}$
@@
4.2. Self-supervised Training of TRADES
@@
Given generation target $\mathbf{x}^{g}_{0}$ and conditional observations $\mathbf{x}^{c}_{0}$, we sample $\mathbf{x}_{t}^{g}=\sqrt{\mkern 2.0mu\overline{\mkern-2.0mu\alpha\mkern-2.0mu}\mkern 2.0mu_{t}}\mathbf{x}_{0}^{g}\,+\,\sqrt{1-\mkern 2.0mu\overline{\mkern-2.0mu\alpha\mkern-2.0mu}\mkern 2.0mu_{t}}\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and train $\boldsymbol{\epsilon}_{\theta}$ by minimizing Eq. (8).
(8) $\mathcal{L}_{\boldsymbol{\epsilon}}(\theta):=\mathbb{E}_{t,\mathbf{x}_{0},\boldsymbol{\epsilon}}\Bigr{[}||\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t}^{g},\mathbf{x}_{0}^{c},t)||^{2}\Bigr{]}.$
@@
Inspired by (Nichol and Dhariwal, 2021), we also learn [^6] Notice from Eq (15) of the original DDPM formulation that $\mathbf{\Sigma}_{\theta}$ is fixed. $\mathbf{\Sigma}_{\theta}$ optimizing it according to Eq. (9).
@@
(9) $\begin{gathered}\mathcal{L}_{\mathbf{\Sigma}}(\theta):=\mathbb{E}_{q}\Biggr{[}\underbrace{-p_{\theta}(\mathbf{x}_{0}^{g}\,|\,\mathbf{x}_{1}^{g},\mathbf{x}_{0}^{c})}_{\text{$L_{0}$}}+\underbrace{D_{KL}(q(\mathbf{x}_{T}^{g}\,|\,\mathbf{x}_{0}^{g})\,||\,p_{\theta}(\mathbf{x}_{T}^{g})}_{\text{$L_{T}$}}\\ +\sum_{t=2}^{T}\underbrace{D_{KL}(q(\mathbf{x}_{t-1}^{g}\,|\,\mathbf{x}_{t}^{g},\mathbf{x}_{0}^{g})\,||\,p_{\theta}(\mathbf{x}^{g}_{t-1}\,|\,\mathbf{x}^{g}_{t},\mathbf{x}_{0}^{c}))}_{\text{$L_{t-1}$}}\Biggr{]}.\end{gathered}$
@@
We optimize Eq. (9) to reduce the negative log-likelihood, especially during the first diffusion steps where $\mathcal{L}_{\Sigma}$ is high (Nichol and Dhariwal, 2021). The final loss function is a linear combination of the two as in Eq: (10).
(10) $\mathcal{L}=\mathcal{L}_{\boldsymbol{\epsilon}}+\lambda\mathcal{L}_{\mathbf{\Sigma}}.$
@@
We perform training by relying on a self-supervised approach as follows. Given a time series $\mathbf{x}_{0}\in\mathbb{R}^{N\times K}$, we apply noise only to its last element – i.e., $\mathbf{x}^{N}_{0}$ – through the forward pass. Then, we denoise it via the reverse process and learn $p_{\theta}(\mathbf{x}^{g}_{t-1}\,|\,\mathbf{x}^{g}_{t},\mathbf{x}_{0}^{c})$ which aims to generate a new sample from the last observations. Therefore, during training, the conditioning has only observed values. A sampling time TRADES generates new samples autoregressively, conditioned on its previous outputs, until the simulation ends.
5. TRADES for Market Simulation
@@
To create realistic and responsive market simulations, we implement TRADES to generate orders conditioned on the market state. TRADES’s objective is to learn to model the distribution of orders. Fig. 1 presents an overview of the diffusion process and the architecture. The network that produces $\varepsilon_{\theta}$ and $\Sigma_{\theta}$ contains several transformer encoder layers to model the temporal and spatial relationships of the financial time series (Sirignano, 2019). Since transformers perform better with large dimension tensors, we project the orders tensor and the LOB snapshots to a higher dimensional space, using two fully connected layers. Hence, TRADES operates on the augmented vector space. After the reverse process, we de-augment $\varepsilon_{\theta}$ and $\Sigma_{\theta}$ projecting them back to the input space to reconstruct $\mathbf{x}_{t-1}^{g}$ and compute the loss.
@@
Conditioning. The conditioned diffusion probabilistic model learns the conditioned probability distribution $p_{\theta}(o|s)$, where $s$ is the market state, and $o$ is the newly generated order, represented as $(p,q,d,b,\delta,c)$, where $p$ is the price, $q$ is the quantity, $d$ is the direction either sell or buy, $b$ is the depth, i.e., the difference between $p$ and the best available price, $\delta$ is the time offset representing the temporal distance from the previously generated order and $c$ is the order type – either market, limit or cancel order. The asymmetries between the buying and selling side of the book indicate shifts in the supply and demand curves caused by exogenous and unobservable factors that influence the price (Cao et al., 2009). Therefore, as depicted in Fig. 1, the model’s conditioning extends beyond the last $N-1$ orders. To effectively capture market supply and demand dynamics encoded within the LOB, we incorporate the last $N$ LOB snapshots of the first $L$ LOB levels as input, where each level has a bid price, bid size, ask price, and ask size. We set $N=256$ as in (Coletta et al., 2021), and $L=10$. We argue that this choice of $L$ is a reasonable trade-off between conditioning complexity and feature informativeness. Several works (Huang and Stoll, 1994; Tran et al., 2022; Pascual and Veredas, 2003; Cao et al., 2008, 2009) have shown that the orders behind the best bid and ask prices play a significant role in price discovery and reveal information about short-term future price movements. In Sec. 7.3, we delve deeper into the conditioning choice and method, performing an ablation and a sensitivity study.
@@
6. DeepMarket framework with synthetic dataset for deep learning market simulations
@@
We present DeepMarket, an open-source Python framework developed for LOB market simulation with deep learning. DeepMarket offers the following features: (1) pre-processing for high-frequency market data; (2) a training environment implemented with PyTorch Lightning; (3) hyperparameter search facilitated with WANDB (Biewald, 2020); (4) TRADES and CGAN implementations and checkpoints to directly generate a market simulation without training; (5) a comprehensive qualitative (via the plots in this paper) and quantitative (via the predictive score) evaluation. To perform the simulation with our world agent and historical data, we extend ABIDES (Byrd et al., 2020), an open-source agent-based interactive Python tool.
6.1. TRADES-LOB: a new synthetic LOB dataset
@@
In LOB research one major problem is the unavailability of a large LOB dataset. In fact, if you want to access a large LOB dataset you need to pay large fees to some data provider. The only two freely available LOB datasets are (Huang et al., 2021) and (Ntakaris et al., [n. d.]) which have a lot of limitations. The first one is composed of only Chinese stocks, which have totally different rules and therefore resulting behaviors with respect to NASDAQ or LSE stocks. The high cost and low availability of lOB data restrict the application and development of deep learning algorithms in the LOB research community. In order to foster collaboration and help the research community we release a synthetic LOB dataset: TRADES-LOB. TRADES-LOB comprises simulated TRADES market data for Tesla and Intel, for two days. Specifically, the dataset is structured into four CSV files, each containing 50 columns. The initial six columns delineate the order features, followed by 40 columns that represent a snapshot of the LOB across the top 10 levels. The concluding four columns provide key financial metrics: mid-price, spread, order volume imbalance, and Volume-Weighted Average Price (VWAP), which can be useful for downstream financial tasks, such as stock price prediction. In total the dataset is composed of 265,986 rows and 13,299,300 cells, which is similar in size to the benchmark FI-2010 dataset (Ntakaris et al., [n. d.]). The dataset will be released with the code in the GitHub repository. We show in the results (section 7.1) how the synthetic market data can be useful to train a deep learning model to predict stock prices.
7. Experiments
@@
Figure 2. PCA analysis on TSLA 29/01. TRADES covers 67.04% of the real distribution, better than the other two methods (52.92% for IABS and 57.49% for CGAN).
@@
Dataset and reproducibility. In almost all SoTA papers in this subfield, the authors use one, two, or three stocks (Shi and Cartlidge, 2023; Coletta et al., 2021; Nagy et al., 2023; Li et al., 2020; Shi and Cartlidge, 2022; Hultin et al., 2023), most of which are tech. Following this practice, we create a LOB dataset from two NASDAQ stocks [^7] The data we used are downloadable from https://lobsterdata.com/tradesquotesandprices upon buying the book indicated on the website, which contains the password to access the data pool. – i.e., Tesla and Intel – from January 2nd to the 30th of 2015. We argue that stylized facts and market microstructure behaviors, which are the main learning objective of TRADES, are independent of single-stock behaviors (see (Bouchaud et al., 2009, 2002; Cont et al., 2014; Gould et al., 2013) [^8] These finance seminal papers discuss universal statistical properties of LOBs across different stocks and markets.), so the particular stock characteristics, such as volatility, market cap, and p/e ratio, are not fundamental. Each stock has 20 order books and 20 message files, one for each trading day per stock, totaling $\sim$24 million samples. The message files contain a comprehensive log of events from which we select market orders, limit orders, and cancel orders. [^9] In LOBSTER, events referring to these orders are defined as deletion.. Each row of the order book file is a tuple $\big{(}P^{ask}(t),V^{ask}(t),P^{bid}(t),V^{bid}(t)\big{)}$ where $P^{ask}(t)$ and $P^{bid}(t)\in\mathbb{R}^{L}$ are the prices of levels $1$ through $L$, and $V^{ask}(t)$ and $V^{bid}(t)\in\mathbb{R}^{L}$ are the corresponding volumes. We use the first 17 days for training, the 18th day for validation, and the last 2 for market simulations. We are aware of the widely used FI-2010 benchmark LOB dataset (Ntakaris et al., [n. d.]) for stock price prediction. However, the absence of message files in this dataset hinders simulating the market since the orders cannot be reconstructed. In Appendix B, we provide an overview of FI-2010 and its limitations.
@@
We conduct the simulations with the first 15 minutes of real orders to compare the generated ones with the market replay. [^10] Market replay denotes the simulation performed with the real historical orders of that day. Afterward, the diffusion model takes full control and generates new orders autoregressively, conditioned on its previous outputs, until the simulation ends. After the world agent generates a new order, there is a post-processing phase in which the output is transformed into a valid order. We begin the simulation at 10:00 and terminate it at 12:00. This choice ensures that the generated orders are sufficient for a thorough evaluation while maintaining manageable processing times. On average, 50,000 orders are produced during this two-hour time frame. The output CSV file of the simulation contains the full list of orders and LOB snapshots of the simulation. All experiments are performed with an RTX 3090 and a portion of an A100. In Appendix C, we detail the data pre- and post-processing and model hyperparameter choice.
@@
Baselines. We compare TRADES with the Market Replay – i.e., ground truth (market replay) – a IABS configuration, and the Wasserstein GAN – i.e., CGAN – under the setting of (Coletta et al., 2022), similar to the same of those proposed in (Coletta et al., 2021, 2023). We implemented CGAN from scratch given that none of the implementations in (Coletta et al., 2023, 2022, 2021) are available. We report details in the Appendix C. Regarding IABS configuration, we used the Reference Market Simulation Configuration, introduced in (Byrd et al., 2020), which is widely used as comparison (Coletta et al., 2021, 2022; Vyetrenko et al., 2020). The configuration includes 5000 noise, 100 value, 25 momentum agents, and 1 market maker.111111the full specifics are in the GitHub page of ABIDES in the tutorial section. We do not compare with other SoTA methods due to the unavailability of open-source implementations and insufficient details to reproduce the results. In some cases, the code is provided but the results are not reproducible due to computational constraints.
7.1. Results

### v3

@@ 4.1 Generation with TRADES
$\boldsymbol{\epsilon}_{\theta}:\left(\mathbf{x}^{g}_{t}\in\mathbb{R}^{S\times K},\>\mathbf{x}^{c}_{0}\in\mathbb{R}^{M\times K},\>t\in\mathbb{R}^{K}\right)\rightarrow\boldsymbol{\epsilon}_{t}\in\mathbb{R}^{S\times K}$ (4)
where $M+S=N$. We set M=N1M=N-1 and $S=1$ for our experiments. Using $S>1$ increases the efficiency, but also the task complexity because, in a single generative step, we generate $S$ time series steps. We exploit the parametrization proposed in [ho2020denoising] described in Eq. (5) to estimate the mean term.
$\boldsymbol{\mu}_{\theta}(\mathbf{x}^{g}_{t},\mathbf{x}^{c}_{0},t)=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}^{g}_{t}-\frac{\beta_{t}}{\sqrt{1-\mkern 1.5mu\overline{\mkern-1.5mu\alpha\mkern-1.5mu}\mkern 1.5mu_{t}}}\boldsymbol{\epsilon}_{\theta}(\mathbf{x}^{g}_{t},\mathbf{x}^{c}_{0},t)\right)\;\>$ (5)
$\text{where}\;\;\>\mathbf{x}_{t}^{g}=\sqrt{\mkern 1.5mu\overline{\mkern-1.5mu\alpha\mkern-1.5mu}\mkern 1.5mu_{t}}\mathbf{x}_{0}^{g}\,+\,\sqrt{1-\mkern 1.5mu\overline{\mkern-1.5mu\alpha\mkern-1.5mu}\mkern 1.5mu_{t}}\boldsymbol{\epsilon}\;\;\;\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (6)
We do not rely on a fixed schedule as in [ho2020denoising] regarding the variance term. Inspired by [nichol2021improved], we learn it as in Eq. (7).
$\mathbf{\Sigma}_{\theta}(\mathbf{x}_{t}^{g},\mathbf{x}_{0}^{c},t)=\exp({v\log\beta_{t}+(1-v)\log\tilde{\beta}_{t}}),$ (7)
where $v$ is the neural network output, together with $\epsilon_{\theta}$. Nichol and Dhariwal [nichol2021improved] found that this choice improves the negative log-likelihood, which we try to minimize. After computing $\boldsymbol{\epsilon}_{t}$ and $\boldsymbol{\sigma}_{t}$, we denoise $\mathbf{x}_{t-1}^{g}$ as in Eq (8).
$\mathbf{x}^{g}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}^{g}_{t}-\frac{\beta_{t}}{\sqrt{1-\mkern 1.5mu\overline{\mkern-1.5mu\alpha\mkern-1.5mu}\mkern 1.5mu_{t}}}\boldsymbol{\epsilon}_{\theta}(\mathbf{x}^{g}_{t},\mathbf{x}^{c}_{0},t)\right)+\boldsymbol{\sigma}_{t}\mathbf{z}$ (8)
@@
4.2 Self-supervised Training of TRADES
@@
Given generation target $\mathbf{x}^{g}_{0}$ and conditional observations $\mathbf{x}^{c}_{0}$, we sample $\mathbf{x}_{t}^{g}=\sqrt{\mkern 1.5mu\overline{\mkern-1.5mu\alpha\mkern-1.5mu}\mkern 1.5mu_{t}}\mathbf{x}_{0}^{g}\,+\,\sqrt{1-\mkern 1.5mu\overline{\mkern-1.5mu\alpha\mkern-1.5mu}\mkern 1.5mu_{t}}\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and train $\boldsymbol{\epsilon}_{\theta}$ by minimizing Eq. (9).
$\mathcal{L}_{\boldsymbol{\epsilon}}(\theta):=\mathbb{E}_{t,\mathbf{x}_{0},\boldsymbol{\epsilon}}\Bigr[||\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t}^{g},\mathbf{x}_{0}^{c},t)||^{2}\Bigr].$ (9)
@@
Inspired by [nichol2021improved], we also learn [^5] Notice from Eq (LABEL:eq_app:6) of the original DDPM formulation that $\mathbf{\Sigma}_{\theta}$ is fixed. $\mathbf{\Sigma}_{\theta}$ optimizing it according to Eq. (10).
@@
$\begin{gathered}\mathcal{L}_{\mathbf{\Sigma}}(\theta):=\mathbb{E}_{q}\Biggr[\underbrace{-p_{\theta}(\mathbf{x}_{0}^{g}\,|\,\mathbf{x}_{1}^{g},\mathbf{x}_{0}^{c})}_{\text{$L_{0}$}}+\underbrace{D_{KL}(q(\mathbf{x}_{T}^{g}\,|\,\mathbf{x}_{0}^{g})\,||\,p_{\theta}(\mathbf{x}_{T}^{g})}_{\text{$L_{T}$}}\\ +\sum_{t=2}^{T}\underbrace{D_{KL}(q(\mathbf{x}_{t-1}^{g}\,|\,\mathbf{x}_{t}^{g},\mathbf{x}_{0}^{g})\,||\,p_{\theta}(\mathbf{x}^{g}_{t-1}\,|\,\mathbf{x}^{g}_{t},\mathbf{x}_{0}^{c}))}_{\text{$L_{t-1}$}}\Biggr].\end{gathered}$ (10)
@@
We optimize Eq. (10) to reduce the negative log-likelihood, especially during the first diffusion steps where $\mathcal{L}_{\Sigma}$ is high [nichol2021improved]. The final loss function is a linear combination of the two as in Eq: (11).
$\mathcal{L}=\mathcal{L}_{\boldsymbol{\epsilon}}+\lambda\mathcal{L}_{\mathbf{\Sigma}}.$ (11)
@@
Where $\lambda$ is used to balance the two terms. We perform training by relying on a self-supervised approach as follows. Given a time series $\mathbf{x}_{0}\in\mathbb{R}^{N\times K}$, we apply noise only to its last element – i.e., $\mathbf{x}^{N}_{0}$ – through the forward pass. Then, we denoise it via the reverse process and learn $p_{\theta}(\mathbf{x}^{g}_{t-1}\,|\,\mathbf{x}^{g}_{t},\mathbf{x}_{0}^{c})$ which aims to generate a new sample from the last observations. Therefore, during training, the conditioning has only observed values. At sampling time, TRADES generates new samples autoregressively, conditioned on its previous outputs, until the simulation ends.
5 TRADES for Market Simulation
@@
To create realistic and responsive market simulations, we implement TRADES to generate orders conditioned on the market state. TRADES’s objective is to learn to model the distribution of orders. Fig. 1 presents an overview of the diffusion process and the architecture. The network that produces $\varepsilon_{\theta}$ and $\Sigma_{\theta}$ contains several transformer encoder layers to model the temporal and spatial relationships of the financial time series [sirignano2019deep]. Since transformers perform better with large dimension tensors, we project the orders tensor and the LOB snapshots to a higher dimensional space, using two fully connected layers. Hence, TRADES operates on the augmented vector space. After the reverse process, we de-augment $\varepsilon_{\theta}$ and $\Sigma_{\theta}$ projecting them back to the input space to reconstruct $\mathbf{x}_{t-1}^{g}$ and compute the loss.
@@
Conditioning. The conditioned diffusion probabilistic model learns the conditioned probability distribution $p_{\theta}(o|s)$, where $s$ is the market state, and $o$ is the newly generated order, represented as $(p,q,d,b,\delta,c)$, where $p$ is the price, $q$ is the quantity, $d$ is the direction either sell or buy, $b$ is the depth, i.e., the difference between $p$ and the best available price, $\delta$ is the time offset representing the temporal distance from the previously generated order and $c$ is the order type – either market, limit or cancel order. The asymmetries between the buying and selling side of the book indicate shifts in the supply and demand curves caused by exogenous and unobservable factors that influence the price [cao2009information]. Therefore, as depicted in Fig. 1, the model’s conditioning extends beyond the last N1N-1 orders. To effectively capture market supply and demand dynamics encoded within the LOB, we incorporate the last $N$ LOB snapshots of the first $L$ LOB levels as input, where each level has a bid price, bid size, ask price, and ask size. We set $N=256$ as in [coletta2021towards], and $L=10$. We argue that this choice of $L$ is a reasonable trade-off between conditioning complexity and feature informativeness. Several works [huang1994market, tran2022informative, pascual2003pieces, cao2008order, cao2009information] have shown that the orders behind the best bid and ask prices play a significant role in price discovery and reveal information about short-term future price movements. In Sec. 7.3, we delve deeper into the conditioning choice and method, performing an ablation and a sensitivity study.
@@
6 DeepMarket framework with synthetic dataset for deep learning market simulations
@@
We present DeepMarket, an open-source Python framework developed for LOB market simulation with deep learning. DeepMarket offers the following features: (1) pre-processing for high-frequency market data; (2) a training environment implemented with PyTorch Lightning; (3) hyperparameter search facilitated with WANDB; (4) TRADES and CGAN implementations and checkpoints to directly generate a market simulation without training; (5) a comprehensive qualitative (via the plots in this paper) and quantitative (via the predictive score) evaluation. To perform the simulation with our world agent and historical data, we extend ABIDES [byrd2020abides], an open-source agent-based interactive Python tool.
6.1 TRADES-LOB: a new synthetic LOB dataset
@@
In LOB research, one major problem is the unavailability of a large LOB dataset. In fact, if you want to access a large LOB dataset, you need to pay large fees to a data provider. The only two freely available LOB datasets are [huang2021benchmark] and [urn:nbn:fi:csc-kata20170601153214969115], which have a lot of limitations. The high cost and low availability of LOB data restrict the application and development of deep learning algorithms in the LOB research community. In order to foster collaboration and help the research community, we release a synthetic LOB dataset: TRADES-LOB. TRADES-LOB comprises simulated TRADES market data for Tesla and Intel for two days. Specifically, the dataset is structured into four CSV files, each containing 50 columns. The initial six columns delineate the order features, followed by 40 columns that represent a snapshot of the LOB across the top 10 levels. The concluding four columns provide key financial metrics: mid-price, spread, order volume imbalance, and Volume-Weighted Average Price (VWAP), which can be useful for downstream financial tasks, such as stock price prediction. In total, the dataset is composed of 265,986 rows and 13,299,300 cells, which is similar in size to the benchmark FI-2010 dataset [urn:nbn:fi:csc-kata20170601153214969115]. The dataset will be released with the code in the GitHub repository. We show in the results (section 7.1) how the synthetic market data can be useful to train a deep learning model to predict stock prices.
7 Experiments
@@
Figure 2: PCA analysis on TSLA 29/01. TRADES covers 67.04% of the real distribution, better than the other two methods (52.92% for IABS and 57.49% for CGAN).
@@
Dataset and reproducibility. In almost all SoTA papers in this subfield, the authors use one, two, or three stocks [shi2023neural, coletta2021towards, nagy2023generative, li2020generating, shi2022state, hultin2023generative], most of which are tech. Following this practice, we create a LOB dataset from two NASDAQ stocks [^6] The data we used are downloadable from https://lobsterdata.com/tradesquotesandprices upon buying the book indicated on the website. – i.e., Tesla and Intel – from January 2nd to the 30th of 2015. We argue that stylized facts and market microstructure behaviors, which are the main learning objective of TRADES, are independent of single-stock behaviors (see [bouchaud2009markets, bouchaud2002statistical, cont2014price, gould2013limit]), so the particular stock characteristics, such as volatility, market cap, and p/e ratio, are not fundamental. Each stock has 20 order books and 20 message files, one for each trading day per stock, totaling $\sim$24 million samples. The message files contain a comprehensive log of events from which we select market orders, limit orders, and cancel orders. Each row of the order book file is a tuple $\big(P^{ask}(t),V^{ask}(t),P^{bid}(t),V^{bid}(t)\big)$ where $P^{ask}(t)$ and $P^{bid}(t)\in\mathbb{R}^{L}$ are the prices of levels $1$ through $L$, and $V^{ask}(t)$ and $V^{bid}(t)\in\mathbb{R}^{L}$ are the corresponding volumes. We use the first 17 days for training, the 18th day for validation, and the last 2 for market simulations. We are aware of the widely used FI-2010 benchmark LOB dataset [urn:nbn:fi:csc-kata20170601153214969115] for stock price prediction. However, the absence of message files in this dataset hinders simulating the market since the orders cannot be reconstructed. In the Appendix [bertiecaiappendix], we provide an overview of FI-2010 and its limitations.
@@
We conduct the simulations with the first 15 minutes of real orders to compare the generated ones with the market replay. [^7] Market replay denotes the simulation performed with the real historical orders of that day. Afterward, the diffusion model takes full control and generates new orders autoregressively, conditioned on its previous outputs, until the simulation ends. After the world agent generates a new order, there is a post-processing phase in which the output is transformed into a valid order. We begin the simulation at 10:00 and terminate it at 12:00. This choice ensures that the generated orders are sufficient for a thorough evaluation while maintaining manageable processing times. On average, 50,000 orders are produced during this two-hour time frame. The output CSV file of the simulation contains the full list of orders and LOB snapshots of the simulation. All experiments are performed with an RTX 3090 and a portion of an A100. In the Appendix [bertiecaiappendix], we detail the data pre- and post-processing and model hyperparameter choice.
@@
Baselines. We compare TRADES with the Market Replay – i.e., ground truth (market replay) – a IABS configuration, and the Wasserstein GAN – i.e., CGAN – under the setting of [coletta2022learning], similar to the same of those proposed in [coletta2021towards, coletta2023conditional]. We implemented CGAN from scratch given that none of the implementations in [coletta2023conditional, coletta2022learning, coletta2021towards] are available. We report details in the Appendix [bertiecaiappendix]. Regarding IABS configuration, we used the Reference Market Simulation Configuration, introduced in [byrd2020abides], which is widely used as comparison [coletta2021towards, coletta2022learning, vyetrenko2020get]. The configuration includes 5000 noise, 100 value, 25 momentum agents, and 1 market maker.888the full specifics are in the GitHub page of ABIDES in the tutorial section. We do not compare with other SoTA methods due to the unavailability of open-source implementations and insufficient details to reproduce the results. In some cases, the code is provided but the results are not reproducible due to computational constraints.
7.1 Results